# DistRL: An Asynchronous Distributed Reinforcement Learning Framework for On-Device Control Agents

### v2

@@ 1 Introduction
The integration of Large Language Models (LLMs) into agents capable of complex tasks has gained momentum with initiatives like AutoGPT Yang et al. (2023a), HuggingGPT Shen et al. (2024), and MetaGPT Hong et al. (2023), AutoUI Zhan & Zhang (2023), etc. These LLM-based agents extend beyond language processing to perform sophisticated functions, including software development and gaming, leveraging their reasoning abilities to interact with and manipulate environments effectively.
@@
One of the key factors driving this trend is the advent of Multimodal Large Language Models (MLLMs), which can process diverse inputs such as text, images, audio, and video, thereby significantly expanding the scope of LLM applications Alayrac et al. (2022), Achiam et al. (2023), Zheng et al. (2024), Li et al. (2023). This versatility enables MLLM-based on-device control agents—intelligent systems embedded within mobile devices that manage and operate applications to execute user commands seamlessly—to interact more naturally and efficiently with their surroundings, completing more complex tasks that require a deeper understanding of context and the ability to learn from interactions. For instance, agents designed to operate smartphone applications can interpret screenshots from the operating system, demonstrating flexibility and adaptability that make them valuable tools in a wide range of scenarios Yang et al. (2023b), Wang et al. (2024). These agents are essential for tasks such as automating app interactions, managing settings, and enhancing user productivity by providing intuitive control over device functionalities.
@@
However, a gap remains between LLMs’ general reasoning capabilities and their effectiveness in GUI-based device control. While LLMs can process information, they struggle with rational behavior and error recovery in real-world environments. Previous solutions use complex wrappers, but without updating model weights, their performance remains limited Bai et al. (2022a). Consequently, prior work in building device agents often relies on constructing complex wrappers around these models, combining them with prompting, search, or tool use. However, without updating the model’s weights, the effectiveness of these agents remains inherently limited by the capabilities of the base model.
@@
To bridge this gap, fine-tuning(type of light training) aforementioned agents on demonstrations via imitation learning Zhan & Zhang (2023), Nakano et al. (2021), Yang et al. (2023b) channels pre-trained abilities into actionable behaviors suitable for device control. However, the dynamic nature of devices and the web renders models trained on static data sub-optimal as ecosystems evolve. Such agents struggle to recover from mistakes in changing environments, limiting their practical utility. Therefore, building robust and reliable device-control agents necessitates an interactive approach that enables MLLMs to adapt and learn from their own experiences on devices and the Internet.
@@
Moreover, MLLMs are prone to confidently generating incorrect content that deviates from human preferences Ouyang et al. (2022), Ziegler et al. (2019), Bai et al. (2022a), Stiennon et al. (2020). To address these challenges, introducing reinforcement learning (RL)-based fine-tuning methods, such as Reinforcement Learning from Human Feedback (RLHF), becomes essential. RLHF leverages human feedback to align model outputs with desired behaviors and preferences. By incorporating RL-based methods, models can learn to optimize policies that not only perform tasks effectively but also adhere to human expectations. Details of the problem formulation as an RL-based fine-tuning approach will be discussed in Sections 2.2 and 3.
@@
One significant challenge in RL-based fine-tuning of mobile agents is the lack of support for efficient online fine-tuning. Existing offline datasets, such as AitW Rawles et al. (2024b) and AndroidControl Li et al. (2024), provide static data that do not capture the dynamic and evolving nature of mobile apps. Frequent updates and new elements like advertisements cause distribution shifts that offline-trained agents struggle to handle, leading to failures in real-world deployments.
@@
To further demonstrate these challenges and the non-trivial nature of bringing MLLMs into mobile device control scenarios, our extensive case studies reveal that even the most advanced proprietary Multimodal Large Language Models (MLLMs) like GPT-4V Achiam et al. (2023), agents such as AutoUI with Supervised Fine-Tuning (SFT), and state-of-the-art mobile control agents like DigiRL Bai et al. (2024) fail in numerous scenarios. A detailed analysis of these failure modes is provided in Appendix A.1.
@@ 2.1 Multi-Modal On-device Control Agents
Recent advancements in pre-trained Large Language Models (LLMs) and Multimodal LLMs (MLLMs) have revolutionized on-device control agents, moving beyond early methods like behavioral cloning or reinforcement learning Osa et al. (2018), Mnih et al. (2015), Shi et al. (2017). Early agents simulated mouse clicks and typing Shi et al. (2017), Humphreys et al. (2022) but faced scalability and adaptability challenges.
Modern approaches use pre-trained models with zero or few-shot prompting and fine-tuning for enhanced capabilities. WebGPT Nakano et al. (2021) employ fine-tuned models for web browsing, while WebAgent Gur et al. (2023) generates web code using T5. AppAgent Yang et al. (2023b) and MobileAgent Wang et al. (2024) act as drivers, enabling the LLMs to explore and act on mobile device environments. Training multimodal device control agents poses challenges like pixel-level interactions and variability in device ecosystems. Many rely on proprietary Vision-Language Models (VLMs) and wrappers for GUI visual grounding Driess et al. (2023), Reid et al. (2024), but without fine-tuning, they are limited by the base models Driess et al. (2023).
Some works fine-tune VLMs with demonstration data, e.g., AutoUI, CogAgent Kapoor et al. (2024), Zhan & Zhang (2023), but static data-trained models may struggle with real-world variability Jiang et al. (2023). Others use filtered imitation learning with autonomously collected data Pan et al. (2024), Lai et al. (2024), While DigiRL Bai et al. (2024) supports on-device RL fine-tuning, it encounters significant inefficiencies in parallel environments. Specifically, DigiRL’s multi-machine setup relies on a fully synchronous data acquisition process, causing faster workers to idle while waiting for slower ones. This approach is impractical in real-world scenarios where task durations can vary by up to 100 times, ranging from seconds to over ten minutes. Additionally, DigiRL lacks support for efficient distributed learning algorithm designs, further hindering its scalability and performance in dynamic, parallel settings. To address this, our scalable and asynchronous RL fine-tuning pipeline, DistRL, offers an efficient solution for distributed mobile control agent training.
@@ 2.2 Reinforcement Learning for On-device Agent Fine-tuning
Reinforcement Learning from Human Feedback (RLHF) is widely used to fine-tune LLMs to align with human preferences Stiennon et al. (2020), Ouyang et al. (2022). In device control tasks, similar approaches use imitation learning from human-labeled evaluations, but RLHF is labor-intensive due to the need for human annotations Ouyang et al. (2022), Bai et al. (2022a).
Recent advances in MLLMs Alayrac et al. (2022), Driess et al. (2023), Reed et al. (2022), Li et al. (2023), Zhang et al. (2023) show impressive multimodal capabilities but often produce incorrect outputs that deviate from human preferences Ouyang et al. (2022), Ziegler et al. (2019), Bai et al. (2022a), Stiennon et al. (2020). Reinforcement Learning from AI Feedback (RLAIF), using AI labelers as proxies, offers an alternative Bai et al. (2022b), Yu et al. (2024), Lee et al. (2023). For on-device tasks, AI evaluators assess task completion using prompts and screenshots Bai et al. (2024), Lee et al. (2023), Yu et al. (2024).
Previous RL research focused on single-turn tasks, limiting their effectiveness for multi-step problems Reed et al. (2022), Liang et al. (2023), Driess et al. (2023). To address this, we developed a simplified off-policy multi-turn RL algorithm, A-RIDE, which learns from suboptimal online interactions, reducing complexity and accelerating convergence compared to previous value-based methods Driess et al. (2023), Nakano et al. (2021), Reed et al. (2022), Yao et al. (2023). This approach is effective for large-scale applications like Android device control.
@@ 2.3 Scalable and Distributed RL Framework
Scalable reinforcement learning frameworks like Ray RLlib Liang et al. (2018) enable distributed training by parallelizing policy learning across CPUs and GPUs using the Ray engine Liang et al. (2017). Ray RLlib supports various algorithms and efficiently manages neural network training, but it assumes that data collection can be simulated or parallelized within the same infrastructure. This assumption limits its applicability to real-world on-device control tasks involving actual mobile devices, where data collection must occur across distributed, heterogeneous devices with varying task durations and network conditions.
Moreoever, frameworks such as IMPALA Espeholt et al. (2018) and IMPACT Luo et al. (2019) are optimized for fast simulations, enabling efficient data collection and policy updates in highly parallelized environments. Applying these frameworks to on-device control introduces non-trivial challenges due to the stochastic nature of real-world interactions and the scalability constraints of mobile devices, including heterogeneous device capabilities, variable task execution times, and unreliable communication.
Our approach builds upon the foundational ideas of IMPALA and IMPACT, effectively extending their concepts to accommodate the unique challenges of on-device control. We introduce a decoupled and asynchronous architecture that supports distributed data collection from heterogeneous, real-world devices by implementing sophisticated communication protocols and optimized data utilization strategies. Furthermore, we develop a novel reinforcement learning algorithm specifically designed to handle the stochasticity and scalability inherent in mobile device environments. This algorithm enhances and maximizes the utilization of collected data while maintaining the exploration behavior of workers, addressing critical aspects that prior work utilizing A3C Mnih (2016) have largely overlooked.
@@ 3 Problem Setup and Preliminaries
A critical component of our RL framework is the ability to obtain reliable reward signals in real-time. To achieve this, we utilize Gemini-1.5-pro Reid et al. (2024) as an autonomous evaluator to assess whether the agent has successfully completed the task at each state. The evaluator receives the current observation, composed of the task description and a screenshot of the device, and outputs a reward signal. Specifically, the evaluator assigns a reward $r_{t}=1$ if the screenshot indicates successful task completion and $r_{t}=0$ otherwise. This effectively transforms the problem into a Partially Observable MDP (POMDP), where the evaluator helps determine the termination condition based on the agent’s observation. Additionally, we applied a reward penalty on unexpected behaviors like repetition which we observed many times in collected trajectories. Details of how we implemented the auto-evaluation can be found in Appendix A.2.
@@ 5 Methodology
Reinforcement learning in distributed device control environments encounters significant challenges related to policy stability, convergence, and effective exploration. Stable convergence is essential for reliable agent performance, while robust exploratory behavior is crucial for discovering effective control policies in dynamic and complex settings. On-policy algorithms such as Proximal Policy Optimization (PPO) Schulman et al. (2017) and Advantage Actor-Critic (A2C) Mnih (2016) are limited by their reliance on synchronous data collection and policy updates, leading to sample inefficiency and delayed learning in asynchronous, multi-agent environments.
@@ 5.1 A-RIDE: The Backbone of DistRL
Our method employs advantage-based estimations to refine policy gradient updates, as an extension of Generalized Advantage Estimation (GAE) Schulman et al. (2015), effectively balancing exploration and exploitation in the learning process. By introducing a trace decay parameter, A-RIDE manages the bias-variance trade-off in advantage calculations, optimizing the stability and convergence of the policy. A-RIDE incorporates enhancements tailored to distributed, asynchronous environments, ensuring robust policy stability and efficient learning in complex device control tasks.
While GAE has proven highly effective in synchronous environments, its reliance on synchronized data collection makes it less suitable for asynchronous settings, such as distributed control tasks on devices. To handle asynchronous trajectory generation in device control tasks, A-RIDE leverages an enhanced Retrace($\lambda$) method for robust off-policy corrections inspired by Espeholt et al. (2018), Munos et al. (2016). Unlike traditional on-policy algorithms that require synchronous data collection, A-RIDE is designed for distributed environments with asynchronous data.
@@
To ensure effective exploration within the action space and prevent the generation of nonsensical or invalid commands, we incorporate entropy regularization into the actor loss function. This addresses the challenge inherent in Vision-Language Models (VLMs) where purely random exploration may lead to semantically incoherent actions Li et al. (2017). The actor loss is defined as:
@@
where the advantage function $A(s_{t},a_{t})$ represents how much better an action $a_{t}$ is compared to the expected reward at state $s_{t}$, given the policy’s understanding of the environment. It can be expressed as the difference between the action-value function $Q(s_{t},a_{t})$ and the state-value function $V(s_{t})$: $A(s_{t},a_{t})=Q(s_{t},a_{t})-V(s_{t})$. The action-value function $Q(s_{t},a_{t})$ estimates the expected return when taking action $a_{t}$ at state $s_{t}$, while the state-value function $V(s_{t})$ estimates the average expected return from state $s_{t}$ under the policy $\pi$. $\mathbb{H}$ is the entropy term, $\mathcal{P}_{\text{invalid}}(a_{t})$ imposes a penalty on actions deemed invalid based on task-specific criteria, $\beta$ controls the strength of entropy regularization, and $\lambda$ modulates the penalty’s influence. The penalty is assigned using validation through pre-trained LLMs like Gemini Reid et al. (2024)), ensuring that only contextually appropriate actions are penalized. The hyperparameters $\beta$ and $\lambda$ are optimized through empirical studies to balance exploration and policy robustness effectively. This formulation encourages the agent to explore a diverse set of actions while constraining it to generate valid and meaningful commands, thereby enhancing both exploration and policy robustness.

### v3

@@ 1 Introduction
The integration of Large Language Models (LLMs) into agents capable of complex tasks has gained momentum with initiatives like AutoGPT (Yang et al., 2023a), HuggingGPT (Shen et al., 2024), and MetaGPT (Hong et al., 2023), AutoUI (Zhan & Zhang, 2023), etc. These LLM-based agents extend beyond language processing to perform sophisticated functions, including software development and gaming, leveraging their reasoning abilities to interact with and manipulate environments effectively.
@@
One of the key factors driving this trend is the advent of Multimodal Large Language Models (MLLMs), which can process diverse inputs such as text, images, audio, and video, thereby significantly expanding the scope of LLM applications (Alayrac et al., 2022, Achiam et al., 2023, Zheng et al., 2024, Li et al., 2023). This versatility enables MLLM-based on-device control agents—intelligent systems embedded within mobile devices that manage and operate applications to execute user commands seamlessly—to interact more naturally and efficiently with their surroundings, completing more complex tasks that require a deeper understanding of context and the ability to learn from interactions. For instance, agents designed to operate smartphone applications can interpret screenshots from the operating system, demonstrating flexibility and adaptability that make them valuable tools in a wide range of scenarios (Yang et al., 2023b, Wang et al., 2024). These agents are essential for tasks such as automating app interactions, managing settings, and enhancing user productivity by providing intuitive control over device functionalities.
@@
However, a gap remains between LLMs’ general reasoning capabilities and their effectiveness in GUI-based device control. While LLMs can process information, they struggle with rational behavior and error recovery in real-world environments. Previous solutions use complex wrappers, but without updating model weights, their performance remains limited (Bai et al., 2022a). Consequently, prior work in building device agents often relies on constructing complex wrappers around these models, combining them with prompting, search, or tool use. However, without updating the model’s weights, the effectiveness of these agents remains inherently limited by the capabilities of the base model.
@@
To bridge this gap, fine-tuning(type of light training) aforementioned agents on demonstrations via imitation learning (Zhan & Zhang, 2023, Nakano et al., 2021, Yang et al., 2023b) channels pre-trained abilities into actionable behaviors suitable for device control. However, the dynamic nature of devices and the web renders models trained on static data sub-optimal as ecosystems evolve. Such agents struggle to recover from mistakes in changing environments, limiting their practical utility. Therefore, building robust and reliable device-control agents necessitates an interactive approach that enables MLLMs to adapt and learn from their own experiences on devices and the Internet.
@@
Moreover, MLLMs are prone to confidently generating incorrect content that deviates from human preferences (Ouyang et al., 2022, Ziegler et al., 2019, Bai et al., 2022a, Stiennon et al., 2020). To address these challenges, introducing reinforcement learning (RL)-based fine-tuning methods, such as Reinforcement Learning from Human Feedback (RLHF), becomes essential. RLHF leverages human feedback to align model outputs with desired behaviors and preferences. By incorporating RL-based methods, models can learn to optimize policies that not only perform tasks effectively but also adhere to human expectations. Details of the problem formulation as an RL-based fine-tuning approach will be discussed in Sections 2.2 and 3.
@@
One significant challenge in RL-based fine-tuning of mobile agents is the lack of support for efficient online fine-tuning. Existing offline datasets, such as AitW (Rawles et al., 2024b) and AndroidControl (Li et al., 2024a), provide static data that do not capture the dynamic and evolving nature of mobile apps. Frequent updates and new elements like advertisements cause distribution shifts that offline-trained agents struggle to handle, leading to failures in real-world deployments.
@@
To further demonstrate these challenges and the non-trivial nature of bringing MLLMs into mobile device control scenarios, our extensive case studies reveal that even the most advanced proprietary Multimodal Large Language Models (MLLMs) like GPT-4V (Achiam et al., 2023), agents such as AutoUI with Supervised Fine-Tuning (SFT), and state-of-the-art mobile control agents like DigiRL (Bai et al., 2024) fail in numerous scenarios. A detailed analysis of these failure modes is provided in Appendix A.1.
@@ 2.1 Multi-Modal On-device Control Agents
Recent advancements in pre-trained Large Language Models (LLMs) and Multimodal LLMs (MLLMs) have revolutionized on-device control agents, moving beyond early methods like behavioral cloning or reinforcement learning (Osa et al., 2018, Mnih et al., 2015, Shi et al., 2017). Early agents simulated mouse clicks and typing (Shi et al., 2017, Humphreys et al., 2022) but faced scalability and adaptability challenges.
Modern approaches use pre-trained models with zero or few-shot prompting and fine-tuning for enhanced capabilities. WebGPT (Nakano et al., 2021) employ fine-tuned models for web browsing, while WebAgent (Gur et al., 2023) generates web code using T5. AppAgent (Yang et al., 2023b) and MobileAgent (Wang et al., 2024) act as drivers, enabling the LLMs to explore and act on mobile device environments. Training multimodal device control agents poses challenges like pixel-level interactions and variability in device ecosystems. Many rely on proprietary Vision-Language Models (VLMs) and wrappers for GUI visual grounding (Driess et al., 2023, Reid et al., 2024), but without fine-tuning, they are limited by the base models (Driess et al., 2023).
Some works fine-tune VLMs with demonstration data, e.g., AutoUI, CogAgent (Kapoor et al., 2024, Zhan & Zhang, 2023), but static data-trained models may struggle with real-world variability (Jiang et al., 2023). Others use filtered imitation learning with autonomously collected data (Pan et al., 2024, Lai et al., 2024), While DigiRL (Bai et al., 2024) supports on-device RL fine-tuning, it encounters significant inefficiencies in parallel environments. Specifically, DigiRL’s multi-machine setup relies on a fully synchronous data acquisition process, causing faster workers to idle while waiting for slower ones. This approach is impractical in real-world scenarios where task durations can vary by up to 100 times, ranging from seconds to over ten minutes. Additionally, DigiRL lacks support for efficient distributed learning algorithm designs, further hindering its scalability and performance in dynamic, parallel settings. To address this, our scalable and asynchronous RL fine-tuning pipeline, DistRL, offers an efficient solution for distributed mobile control agent training.
@@ 2.2 Reinforcement Learning for On-device Agent Fine-tuning
Reinforcement Learning from Human Feedback (RLHF) is widely used to fine-tune LLMs to align with human preferences (Stiennon et al., 2020, Ouyang et al., 2022). In device control tasks, similar approaches use imitation learning from human-labeled evaluations, but RLHF is labor-intensive due to the need for human annotations (Ouyang et al., 2022, Bai et al., 2022a).
Recent advances in MLLMs (Alayrac et al., 2022, Driess et al., 2023, Reed et al., 2022, Li et al., 2023, Zhang et al., 2023) show impressive multimodal capabilities but often produce incorrect outputs that deviate from human preferences (Ouyang et al., 2022, Ziegler et al., 2019, Bai et al., 2022a, Stiennon et al., 2020). Reinforcement Learning from AI Feedback (RLAIF), using AI labelers as proxies, offers an alternative (Bai et al., 2022b, Yu et al., 2024, Lee et al., 2023). For on-device tasks, AI evaluators assess task completion using prompts and screenshots (Bai et al., 2024, Lee et al., 2023, Yu et al., 2024).
Previous RL research focused on single-turn tasks, limiting their effectiveness for multi-step problems (Reed et al., 2022, Liang et al., 2023, Driess et al., 2023). To address this, we developed a simplified off-policy multi-turn RL algorithm, A-RIDE, which learns from suboptimal online interactions, reducing complexity and accelerating convergence compared to previous value-based methods (Driess et al., 2023, Nakano et al., 2021, Reed et al., 2022, Yao et al., 2023). This approach is effective for large-scale applications like Android device control.
@@ 2.3 Scalable and Distributed RL Framework
Scalable reinforcement learning frameworks like Ray RLlib (Liang et al., 2018) enable distributed training by parallelizing policy learning across CPUs and GPUs using the Ray engine (Liang et al., 2017). Ray RLlib supports various algorithms and efficiently manages neural network training, but it assumes that data collection can be simulated or parallelized within the same infrastructure. This assumption limits its applicability to real-world on-device control tasks involving actual mobile devices, where data collection must occur across distributed, heterogeneous devices with varying task durations and network conditions.
Moreoever, frameworks such as IMPALA (Espeholt et al., 2018) and IMPACT (Luo et al., 2019) are optimized for fast simulations, enabling efficient data collection and policy updates in highly parallelized environments. Applying these frameworks to on-device control introduces non-trivial challenges due to the stochastic nature of real-world interactions and the scalability constraints of mobile devices, including heterogeneous device capabilities, variable task execution times, and unreliable communication.
Our approach builds upon the foundational ideas of IMPALA and IMPACT, effectively extending their concepts to accommodate the unique challenges of on-device control. We introduce a decoupled and asynchronous architecture that supports distributed data collection from heterogeneous, real-world devices by implementing sophisticated communication protocols and optimized data utilization strategies. Furthermore, we develop a novel reinforcement learning algorithm specifically designed to handle the stochasticity and scalability inherent in mobile device environments. This algorithm enhances and maximizes the utilization of collected data while maintaining the exploration behavior of workers, addressing critical aspects that prior work utilizing A3C (Mnih, 2016) have largely overlooked.
@@ 3 Problem Setup and Preliminaries
A critical component of our RL framework is the ability to obtain reliable reward signals in real-time. To achieve this, we utilize Gemini-1.5-pro (Reid et al., 2024) as an autonomous evaluator to assess whether the agent has successfully completed the task at each state. The evaluator receives the current observation, composed of the task description and a screenshot of the device, and outputs a reward signal. Specifically, the evaluator assigns a reward $r_{t}=1$ if the screenshot indicates successful task completion and $r_{t}=0$ otherwise. This effectively transforms the problem into a Partially Observable MDP (POMDP), where the evaluator helps determine the termination condition based on the agent’s observation. Additionally, we applied a reward penalty on unexpected behaviors like repetition which we observed many times in collected trajectories. Details of how we implemented the auto-evaluation can be found in Appendix A.2.
@@ 5 Methodology
Reinforcement learning in distributed device control environments encounters significant challenges related to policy stability, convergence, and effective exploration. Stable convergence is essential for reliable agent performance, while robust exploratory behavior is crucial for discovering effective control policies in dynamic and complex settings. On-policy algorithms such as Proximal Policy Optimization (PPO) (Schulman et al., 2017) and Advantage Actor-Critic (A2C) (Mnih, 2016) are limited by their reliance on synchronous data collection and policy updates, leading to sample inefficiency and delayed learning in asynchronous, multi-agent environments.
@@ 5.1 A-RIDE: The Backbone of DistRL
Our method employs advantage-based estimations to refine policy gradient updates, as an extension of Generalized Advantage Estimation (GAE) (Schulman et al., 2015), effectively balancing exploration and exploitation in the learning process. By introducing a trace decay parameter, A-RIDE manages the bias-variance trade-off in advantage calculations, optimizing the stability and convergence of the policy. A-RIDE incorporates enhancements tailored to distributed, asynchronous environments, ensuring robust policy stability and efficient learning in complex device control tasks.
While GAE has proven highly effective in synchronous environments, its reliance on synchronized data collection makes it less suitable for asynchronous settings, such as distributed control tasks on devices. To handle asynchronous trajectory generation in device control tasks, A-RIDE leverages an enhanced Retrace($\lambda$) method for robust off-policy corrections inspired by (Espeholt et al., 2018, Munos et al., 2016). Unlike traditional on-policy algorithms that require synchronous data collection, A-RIDE is designed for distributed environments with asynchronous data.
@@
To ensure effective exploration within the action space and prevent the generation of nonsensical or invalid commands, we incorporate entropy regularization into the actor loss function. This addresses the challenge inherent in Vision-Language Models (VLMs) where purely random exploration may lead to semantically incoherent actions (Li et al., 2017). The actor loss is defined as:
@@
where the advantage function $A(s_{t},a_{t})$ represents how much better an action $a_{t}$ is compared to the expected reward at state $s_{t}$, given the policy’s understanding of the environment. It can be expressed as the difference between the action-value function $Q(s_{t},a_{t})$ and the state-value function $V(s_{t})$: $A(s_{t},a_{t})=Q(s_{t},a_{t})-V(s_{t})$. The action-value function $Q(s_{t},a_{t})$ estimates the expected return when taking action $a_{t}$ at state $s_{t}$, while the state-value function $V(s_{t})$ estimates the average expected return from state $s_{t}$ under the policy $\pi$. $\mathbb{H}$ is the entropy term, $\mathcal{P}_{\text{invalid}}(a_{t})$ imposes a penalty on actions deemed invalid based on task-specific criteria, $\beta$ controls the strength of entropy regularization, and $\lambda$ modulates the penalty’s influence. The penalty is assigned using validation through pre-trained LLMs like Gemini (Reid et al., 2024)), ensuring that only contextually appropriate actions are penalized. The hyperparameters $\beta$ and $\lambda$ are optimized through empirical studies to balance exploration and policy robustness effectively. This formulation encourages the agent to explore a diverse set of actions while constraining it to generate valid and meaningful commands, thereby enhancing both exploration and policy robustness.